If the teacher is 37 years old and one of her students is 11 years old, in how many years will she be twice as old as her students?

Found 2 solutions by basketball1127, jgr45:
Answer by basketball1127(19)

Let x be the number of years
37+x = 2(11+x)
37+x = 22+2x
2x-x = 37-22
x = 15 years

Answer by jgr45(31)

Admittedly you can probably get better answers from other tutors on here but this is my mind's process solving this problem:
The teacher is 37, the student is 11. That's a difference of 26 years. So, 26 has to be the age of the student when the teacher is twice that, or 52; and as you can see, student and teacher will reach those ages in 15 years.